OpenStudy (anonymous):

A passenger plane mad a trip to Las Vegas and back. On the trip there it flew 432 mph and on the return trip it went 480 mph. How long did the trip there take if the return trip took nine hours?

OpenStudy (austinl):

I would set up a proportion, do you know how to do this?

OpenStudy (anonymous):

My teacher has explained it, but Im lost

OpenStudy (austinl):

\(\displaystyle \frac{480}{9}=\frac{432}{x}\) Solve for x.

OpenStudy (anonymous):

Ok I got it! thank you

OpenStudy (austinl):

No problemo!